A cuvette with a path length of 1.2 cm and solution containing 0.86 mol/dm3

of a dye

gives an absorbance of 0.8. Calculate the molar extinction coefficient of the dye.


1
Expert's answer
2021-04-21T03:53:52-0400

Solution:

Beer-Lambert Law:

A = ε × l × C

where:

A = absorbance (no units)

ε = molar extintion coefficient (M-1 cm-1)

l = pathlength (cm)

C = concentration (M)


Thus:

ε = A / (l × C)

ε = (0.8) / (1.2 cm × 0.86 M) = 0.775 M-1 cm-1

ε = 0.775 M-1 cm-1


Answer: The molar extinction coefficient (ε) of the dye is 0.775 M-1 cm-1.